cylinder of the anesthetic gas, isoflurane (C3H2ClF5O) in an operating room has a pressure of
7754 torr. The temperature in the room is 25oC, and the cylinder has a volume of 10.00 L. How
many grams of isoflurane are still present in the tank?
Answer #1

Given:
P = 7754.0 torr
= (7754.0/760) atm
= 10.2026 atm
V = 10.0 L
T = 25.0 oC
= (25.0+273) K
= 298 K

find number of moles using:
P * V = n*R*T
10.2026 atm * 10 L = n * 0.08206 atm.L/mol.K * 298 K
n = 4.172 mol

Molar mass of C3H2ClF5O,
MM = 3*MM(C) + 2*MM(H) + 1*MM(Cl) + 5*MM(F) + 1*MM(O)
= 3*12.01 + 2*1.008 + 1*35.45 + 5*19.0 + 1*16.0
= 184.496 g/mol

use:
mass of C3H2ClF5O,
m = number of mol * molar mass
= 4.172 mol * 1.845*10^2 g/mol
= 7.698*10^2 g
Answer: 7.70*10^2 g